Billy Madison consumes 100 units of X and 50 units of Y. The price of x rises from 2 to 3. The price of Y remains at 4.


a/ how much must Billy’s income rise so that he can exactly afford 100 units of x and 50 units of y?


b/ Draw Billy’s original budget set when the price of X is 2


c/ Draw Billy’s budget set when the price of X is 3, but his income is unchanged


d/ Draw Billy’s budget set when the price of X is 3 and his income increases by the amount in (a) so that he can afford his original consumption bundle.


1
Expert's answer
2021-11-10T10:06:48-0500

a.

"income=p_xx+p_yy\\x=100,y=50\\p_x=2,p_y=4\\income=2(100)+4(50)=400"

now px rises to 3 , but x=100,y=50

"income=3(100)+4(50)=500\\"

income must rise by "=500-400=100"


b and c

original budget "400=2x+4y"

x intercept "=\frac{400}{2}=200"

y intercept "=\frac{400}{4}=100"

new budget "400=3x+4y"

x intercept "=\frac{400}{3}=133.33"

y intercept "=\frac{400}{4}=100"



d.

new budget "500=3x+4y"

x intercept "=\frac{500}{3}=166.67"

y intercept "=\frac{500}{4}=125"